OpenStudy (anonymous):

Which pair are solutions to the equation? 4xy + 8 = 36 A. (1, 7) and (7, 1) B. (7, 1) and (3, 2) C. (4, 9) and (3, 2) D. (1, 7) and (4, 9)

OpenStudy (anaise):

cx I'm robin

OpenStudy (anaise):

I killed her

OpenStudy (anonymous):

:D

pooja195 (pooja195):

Just plug your options into the equation (given)

OpenStudy (anonymous):

she's working on a tv show right now shes still alive

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[4xy + 8 = 36\] put \(x=7,y=1\) and get \[4\times 1\times 7+8=28+8=36\]so that one works

pooja195 (pooja195):

Or that .-.

OpenStudy (anaise):

:D

OpenStudy (anonymous):

its confusing

OpenStudy (anaise):

cornfrused***

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 Just plug your options into the equation (given) \(\color{#0cbb34}{\text{End of Quote}}\) This?

OpenStudy (anonymous):

HOLY pellet! 100 SMARTSCORE????????

OpenStudy (anonymous):

for \( (7, 1)\) put \(x=7,y=1\) get \[4\times 7\times 1+8=28+8=36\] looks like we have another winner

OpenStudy (just_one_last_goodbye):

Lol xD nice attention title btw

OpenStudy (anaise):

you never saw the number 100?

OpenStudy (theedwardsfamily):

^^^^

OpenStudy (anonymous):

nope....

OpenStudy (anaise):

what do you have......99 bucks?

OpenStudy (anaise):

65?

OpenStudy (anonymous):

*middle finger to anaise*

OpenStudy (just_one_last_goodbye):

@satellite73 can you leave OS so I can be the new 100 SS? ^_^ I've heard rumors if a user has a 100 SS they have to wait for that person to LEAVE

OpenStudy (anonymous):

https://www.youtube.com/watch?v=vWUn6tzOAwU

pooja195 (pooja195):

Congrats for gathering everyones attention c:

OpenStudy (theedwardsfamily):

xDDD

pooja195 (pooja195):

@satellite73 its so old it makes me cri ;-; ew the music

OpenStudy (samsungfanboy):

Batman is dead.. Satellite ran over him with his bike T_T

OpenStudy (anaise):

Robin is in charge

OpenStudy (anaise):

danny is starfire

OpenStudy (theedwardsfamily):

LOL What about Wonder Woman? ;)

OpenStudy (anonymous):

it's the joker. Quick Hide under my parade

OpenStudy (anonymous):

well how else am i going to get help openstudy was suppose to be s site where you get help and i saw the video and thats what i thought ... wells you do but you have to get redicluslaus

OpenStudy (just_one_last_goodbye):

@satellite73 if you want to make a joke about batman make it funny by making Batman + Taylor swift = this song https://www.youtube.com/watch?v=maX-ObiJB3o

OpenStudy (anaise):

ooh....Im switching characters

OpenStudy (anaise):

I am harley quinn

OpenStudy (anonymous):

I'm smarter than Anaise xD

OpenStudy (theedwardsfamily):

lol guess I'll be Black Widow! ^_^

OpenStudy (anonymous):

noone is smarter then anyone

OpenStudy (anonymous):

I'm smarter than MariahDiamond also

OpenStudy (anonymous):

depending on smartscores

OpenStudy (anaise):

otay jinx e.e

OpenStudy (anonymous):

*okay

OpenStudy (anaise):

says who? you could have been here longer. I just joined yesterday c:

OpenStudy (anonymous):

Yeah right I saw you a week ago and I was joking about Jaynator

OpenStudy (anaise):

haha...you're illusive stalker

OpenStudy (anonymous):

solve 7|x|=42

OpenStudy (anaise):

why may I ask? :)

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

adios harley

OpenStudy (anaise):

I was just asking, if you want a answer I would happily give it

OpenStudy (anonymous):

adios Mariah you're beautiful. (lesbian speaking)

OpenStudy (anaise):

but I owe you nothing, nor an explanation

OpenStudy (anonymous):

thanks ?

OpenStudy (anonymous):

at lest help me anaise

OpenStudy (anaise):

I did, I told you batman was dead...-_-

OpenStudy (anonymous):

NOT THAT THE QUESTION IS BELOW IT

OpenStudy (anaise):

^_^ ask the seri

OpenStudy (anonymous):

....

OpenStudy (anaise):

cx

OpenStudy (mepitt):

i think its a

OpenStudy (anonymous):

FINALLY A PERISON WITH A LIGIT ANSWER

OpenStudy (mepitt):

lol

OpenStudy (mepitt):

welcome;)

alones (alones):

wait so what was the answer??